Shift report 20030214 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Fri Feb 14 2003 - 15:51:39 EST

  • Next message: Apache: "Shift report 20030215 16:00-24:00"
    ------------------------------------------------------------
    Date: Fri Feb 14 15:51:39 2003, Shift: 08:00-16:00
    Supervisor: Eun-Joo Kim, with: 
    Log book#14 pages 16 - 16
    
    Shift Summary:
    ------------------------------------------------------------
    1. We had a long store from 7am to 3pm.
       Stopped taking-data at 3pm, and gap cleaning started
       at 3:30pm. Liminosity was not great. 
       (ZDC rate 2159[run7891] -> 1373[run7893])
       TOFW/ZDC rate was about 4-6. 
    
    2. Data Run :
    
       We finished FS 1/20A at 12-degree setting,
       though mainframe L1458B(BFS & RICH) HV was off.
    
       7890,7891,7892 : MRS 90-degree, 1050B
                        FS  12-degree, 1/20A 
       7893           : MRS 90-degree, 1050B
                        FS  12-degree, 1/8A
    
       DVM run : 7895
    
    3. TD1(U1,U3,D3) & MRST3 HV were re-setting, and changed
       BRAHVO TRT0 default setting.
       H2 slat top7, bottom 15 : need to fix or check?
       tsir terminal server for TOFW rebooted.
    
    Plan :
    
     MRS : keep current setting, 90-degree 1050B
     FS  : need more 180K trigger6 at 1/8A 12-degree setting
      
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Fri Feb 14 2003 - 15:53:50 EST